OpenStudy (pawanyadav):

Physics question :- in an elastic collision -The final linear momentum is equal to initial linear momentum. Is this correct? Why?

OpenStudy (irishboy123):

it follows from Newton's Laws. 3rd law, action and reaction are equal and opposite; and 2nd law \(F = \dot p\). so for a collision where \(F ~dt = d(mv)\), we have \( F_1 = -F_2, \quad d(mv)_1 + d(mv)_2 = 0\) and it's also true in an inelastic collision

OpenStudy (pawanyadav):

So this is correct

OpenStudy (pawanyadav):

But that is not the answer ,it is incorrect

OpenStudy (pawanyadav):

Oh I think the initial linear momentum is negative of final linear momentum and vice versa ......that's why it is incorrect.

ganeshie8 (ganeshie8):

linear momentum of a closed and isolated "system" cannot change when there is no external force

ganeshie8 (ganeshie8):

linear momentum of individual objects can change, but the total linear momentum remains constant

OpenStudy (pawanyadav):

Then answer should be correct but it's not.

ganeshie8 (ganeshie8):

your question is ambiguous

ganeshie8 (ganeshie8):

``` -The final linear momentum is equal to initial linear momentum. Is this correct? Why? ``` final linear momentum of what ?

OpenStudy (pawanyadav):

Obviously, it should be for complete system

ganeshie8 (ganeshie8):

assuming no external force, the linear momentum of system must be same at any point in time

Parth (parthkohli):

The law of conservation of momentum is not applicable for, say, this system.

Parth (parthkohli):

Try to answer why.

OpenStudy (pawanyadav):

If we Take example of an elastic collision between a ball and wall Initial momentum is =MV {M is mass of ball and V velocity) And final =-MV (as wall is fixed so velocity of ball is reversed.)

ganeshie8 (ganeshie8):

the system must be closed and isolated

OpenStudy (pawanyadav):

Because gravity acts as external force

Parth (parthkohli):

Plus there are impulsive adhesive forces in the wall.

Parth (parthkohli):

No, not gravity